# secReferences in the Online Appendix

### v2

@@ 4 Timing Campaigns for Media Exposure
We start our analysis by investigating when NGOs are most likely to launch their campaigns relative to a firm’s corporate calendar. Figure 1(a) shows that NGOs are significantly more likely to launch campaigns on a firm’s AGM date. Campaigns are about 4 to 6 times more common on AGM weeks compared to other weeks; while a typical week sees 28 campaigns, AGM weeks have more than 100 campaigns. [^7] Appendix Figure B3 confirms the result showing a significant drop in campaign density immediately after AGM dates (p-value $\leq 1\%$), following Cattaneo et al. (2020).
@@
(a) Timing of NGO campaigns
(b) Media coverage in the AGM week (gray)
@@
This rise could be related to the heightened media coverage during AGMs (see also Fioretti et al., 2023), as illustrated in Figure 1(b), where news coverage peaks in the 10 days around an AGM (gray bars) compared to other periods (white bars). To test if media exposure of NGO campaigns is highest around AGMs, we aggregate news by firm ($n$) and month-by-year ($m,y$) and run the following event study:
@@
$\text{News}_{n,m,y}$ $=\sum^{\tau=6}_{\begin{subarray}{c}\tau=-6\\ \tau\neq-3\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}_{\{m=\tau\}}\cdot\text{NGO Campaign}_{n,m,y}\right)+\mathbf{X}_{n,m,y}\mathbf{\gamma}^{\prime}+\alpha_{n,y}+\alpha_{m,y,s(n)}+u_{n,m,y}.$ (1)
First, note that we adjusted the calendar year to align with the corporate year so that the year starts six months before the AGM month (month 0) and ends six months later for each firm. Importantly, since firms are unlikely to change their AGM month over time due to SEC regulations as outlined in Section 3, moving from the calendar year to the corporate year does not affect the distribution of months by firms (i.e., months in following years are not double counted). Thus, the recalendarization is a mere relabeling of the time variation within a firm. [^8] In each year, we define months -6 and +6 as half months, so the calendar year has twelve full months.
Second, Equation (1) interacts the event-study dummy, $\mathbf{1}_{\{m=\tau\}}$, with the variable $\text{NGO Campaign}_{n,m,y}$, which is 1 if firm $n$ experienced an NGO campaign in month $m$ of year $y$, and 0 otherwise. We use $\tau=-3$ as the reference period because shareholders typically receive voting materials within 90 days of the AGM, marking a date when the AGM’s visibility in the news could increase (ISS, 2022).
@@
The vector $\mathbf{X}_{n,m,y}$ contains the direct effects of $\mathbf{1}_{\{m=\tau\}}$ and $\text{NGO Campaign}_{n,m,y}$ on the dependent variable. Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-month-by-year fixed effects ($\rho_{m,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs). Standard errors are clustered at the firm level.
Figure 2 reports the estimated coefficients, revealing that media coverage peaks during the month of the AGM. To quantify its magnitude, an NGO campaign launched in the AGM month increases media coverage by 80% compared to a campaign started 6 months earlier, on average. After the AGM, coverage returns to average levels, indicating that NGOs gain more media attention by campaigning close to the AGM date.
@@
Figure 2: Media exposure of NGO campaigns around an AGM
Note: We regress the number of news about firm $f$ in month $t$ on a dummy indicating the publication of a negative NGO report about firm $f$ in month $t$, a dummy equal to one when month $t$ is $k$ months away from the nearest AGM, where we take $k$ to vary between -6 and +6 months, the interaction between these two dummy variables, and firm and month by industry fixed effects. Month -6 and +6 are “half” months, so the year still consists of 12 full months. The figure also reports the 95% confidence interval when the standard errors are clustered at the level of the firm. The sample period goes from January 2010 to December 2019 and covers 4,676 firms. NGO campaign data from Sigwatch. AGM data from ISS. News data from Ravenpack. Industry data from Compustat. The regression also includes firm and year-by-industry fixed effects. Standard errors are clustered at the firm level.
@@ 4.1 Stakeholders’ Reactions
In this section, we explore how campaign timing affects critical stakeholders of the targeted firms, such as shareholders and consumers. Therefore, timing a campaign on an AGM date has significant implications with consequences that go beyond media coverage. We report the exhibits related to the analysis in this section in the Online Appendix.
@@ 4.1.1 Shareholders
Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022). Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest. We define as “related” a shareholder proposal on the same topic of the campaign in the previous year. Appendix C describes the matching procedure.
@@
Figure B4 shows that only NGO campaigns timed twelve months prior to the next AGM increase the likelihood of a shareholder proposal on the same topic at the AGM. This substantial increase is as large as 12 percentage points, or 44%, above the average probability of observing a proposal. Thus, NGO campaigns from the previous AGMs likely lead to a shareholder proposal at the current AGM, potentially due to heightened media attention.
@@ 4.1.2 Consumers
Drawing from quarterly data from Compustat, we investigate the impact of NGO campaigns on consumers as the stakeholder consuming a firm’s values (Conway and Boxell, 2024) by studying how the Sales to Asset ratio changed around the time of an NGO campaign to the AGM date through
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}_{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively. The dependent value rebases $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms). That is, we investigate potential consumer boycotts in the months following an NGO campaign. Standard errors are clustered at the firm level.
Appendix Figure B5 reports the estimated $\beta_{\tau}$ coefficients separately for two subsets of the data: for NGO campaigns (i) not planned on an AGM date (blue dots) and (ii) those planned on an AGM date (red triangles). We find that consumers do not respond to NGO campaigns not timed on an AGM day, but they do so two quarters after a campaign targeted on the AGM day, implying a loss of sales to total assets of 0.7pp, or about 5%. Similar sale losses as a result of negative ESG corporate events have been documented in the literature (Conway and Boxell, 2024; Saint-Jean, 2024).
@@
NGO campaigns on an AGM date impact stakeholders. (i) Shareholders are more likely to make related proposals in future years. (ii) Consumers are more likely to boycott the firm in the following months.
5 Timing Campaigns to Influence AGMs
Early campaigns could benefit an NGO aiming to influence current AGM outcomes. Since most votes are cast by proxy before the AGM, targeting the AGM date may limit an NGO’s ability to sway a significant share of the votes. For instance, over 85% of vote-eligible shares for Microsoft’s 2023 AGM were cast by proxy before the meeting (Microsoft, 2023), suggesting that a campaign close to the AGM date may have less impact on the voting results.
@@ 5 Timing Campaigns to Influence AGMs
where $\text{Success}_{i,n,m,y}$ is a dummy indicating the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+l,t}$ is a dummy indicating the occurrence of an NGO campaign on the same topic as proposal $i$, $l$ months before the AGM. The reference period covers the months that occur more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$. We account for firm-by-year, and topic-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and topics within firm-year pairs. The standard errors are clustered by firm.
Figure 3(a) plots the estimated $\hat{\beta}_{l}$. We find that NGO campaigns happening 6 months and 1 month before the AGM positively influence the outcome of related shareholder proposals. They increase, respectively, the probability of success by 2pp (6%) and 3pp (8%) more than a campaign launched more than seven months before.
To understand why campaigns 6 months before the AGM are also somewhat successful, we estimate (4) on the subset of voted proposals, with the percentage of votes in favor as the dependent variable. Note that every proposal to be voted on at the AGM is accompanied by a vote recommendation by the management (“For”, “Against”, or “Abstain”). Institutional investors can also follow the recommendations of proxy advisors, the most popular one being ISS (Shu, 2024). Those recommendations have a critical influence on the vote decisions. We thus include management and ISS’s recommendations as covariates in (4).
Figure 3(b) shows that only NGO campaigns occurring a month or three months before the AGM increases support for related shareholder proposals. This increase is by nearly 2 percentage points. Given the generally low baseline support for these proposals (24%), this translates to a 8% increase in shareholder support when an NGO campaign addresses the same topic, on average. In contrast to Figure 3(a), there is no significant increase in support for campaigns initiated as early as six months before the AGM. This suggests that the difference between the two figures is due to shareholder proposals being withdrawn after reaching an agreement with the firm.
Figure 3: NGOs’ influence on the vote of E&S shareholder proposals
@@
(b) Shareholder support for proposal
Note: E&S denote environmental and social proposals. The figure plots the estimated coefficients of regressions of a dummy indicating that a shareholder proposal is withdrawn or passes if it goes to a vote (left panel, based on 911 proposals), or the percentage of votes in favor of the proposal when it goes to the vote (right panel, based on 2,078 proposals), on dummies indicating the occurrence of an NGO campaign on the same topic $k$ months before the AGM, for $k$ varying from -6 to 0. The regressions also include month, firm-by-year, and topic-by-year fixed effects. Standard errors are clustered at the firm level.
Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015). For this reason, we investigate all E&S proposals in Figure 3(a). Given the low support of proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote. However, those votes can still be helpful in disciplining firms (Levit and Malenko, 2011) as large shareholder support to a cause may be hard to conceal for managers.
@@
NGO campaigns are more effective at influencing a firm’s decisions when launched shortly before the AGM.
@@
Therefore, Sections 4 and 5 outline a tradeoff: late campaigns impact society by boosting an NGO’s visibility among the media and key stakeholders, while early ones are more effective at influencing a firm’s decisions. How do NGOs navigate this tradeoff?
@@
6.1 Campaign Timing and Accumulated NGO Visibility
Section 4 suggests that targeting a firm’s AGM increases an NGO’s visibility in the media and among stakeholders. This visibility could enhance an NGO’s reputation for commitment to a cause. In this section, we explore whether NGOs, after establishing such a reputation, begin launching campaigns earlier—before the target firm’s AGM—to improve their chances of influencing the outcome as suggested in Section 5.
@@ 6.1 Campaign Timing and Accumulated NGO Visibility
$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}_{\{\text{Campaign \#}_{c(\cdot),g,y}=\tau\}}+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$ (5)
which examines $\text{Distance to AGM}_{c(i,n,g),y}$, the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using a dummy, $\text{Campaign \#}_{c,g,y}$, that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise.
Since our campaign data start in 2010, we do not know how many campaigns an NGO launched before 2010. To control for this truncation as well as time varying confounders—e.g., the size of the NGO social media network across years (Tufekci, 2017)—we include NGO-by-year fixed effects. We also include firm-by-year and topic-by-year fixed effects to account for firm and topic-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign. Hence, the identification of $\beta_{\tau}$ comes from multiple campaigns $c(\cdot)$ by the same NGO $g$ within year $y$ [^9] In Appendix Figure B6, we present similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years. Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample. The standard errors are clustered at the NGO level.
To interpret the estimated $\beta_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM. Thus, Figure 4 indicates that as the NGO’s experience increases, the new campaigns are launched before the AGM date. [^10] The large negative coefficients estimated for NGOs with many campaigns—e.g., more than 9—are driven by outlier NGOs with several campaigns per year. We postpone the interpretation of this result to Section 7, which sums up and discuss the results.
@@
Alternative mechanism—selection. NGO do not increasingly launch campaigns against firms that may be more easily influenced over time. Appendix Figures B8 and B9 exclude selection on observables. For instance, target sizes do not covary (market capitalization) with NGO experience. In addition, environmental (social) scores do not covary with an NGO’s experience in environmental (social) campaigns.
@@
Figure 4: NGO experience and the timing of NGO campaigns
Note: This figure plots the estimated coefficients from the regression of the distance between NGO campaigns and the targeted firm’s next AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. Regressions are based on a sample of 1,136 NGO campaigns. Standard errors are clustered at the NGO level.

### v3

@@ 4 Timing Campaigns for Media Exposure
NGOs are substantially more likely to launch their campaigns around AGM dates, on average. According to Figure 1(a), NGO campaigns are about 4 to 6 times more common on AGM weeks compared to other weeks; while a typical week sees 28 campaigns, AGM weeks have more than 100.666Appendix Figure B3 confirms this finding showing a significant drop in campaign density immediately after AGM dates (p-value $\leq 1\%$) (Cattaneo et al., 2020).
@@
(a) Campaigns peak during AGMs
(b) Media coverage rises during AGMs
@@
This finding could be related to the heightened media coverage during AGMs (Fioretti et al., 2023), as illustrated in Figure 1(b), where news coverage peaks in the 10 days around an AGM (gray bars) compared to other periods (white bars). To test if media exposure of NGO campaigns is highest during the AGM week, we aggregate negative E&S news by firm ($n$) and week-by-year ($w,y$) and perform the event study:
@@
$\text{News}_{n,w,y}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}{\{M_{n,w,y}=\tau\}}\,\cdot\,\text{NGO Campaign}_{n,w,y}\right)$ (1)
$+\mathbf{X}_{n,w,y}\mathbf{\gamma}^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y}\,\,,$
where “$\text{News}_{n,w,y}$” is an indicator for whether firm $n$ in week $w$ of year $y$ experiences at least one negative E&S news event. The independent variables include a set of dummies “$\mathbf{1}\{M_{n,w,y}=\tau\}$” indicating the distance in months from the closest AGM. The function $M_{n,w,y}$ assigns week $w$ to each month; the indicator is 1 if the week falls in month $\tau$, where a month is made up of 4 weeks, and zero otherwise.777 The analysis aligns the calendar year with the corporate year, starting five months before the AGM month (month 0) and ending five months later, with months -5 and +5 slightly longer. Since SEC regulations make AGM months stable over time (Section 3), this relabeling does not alter the distribution of months by firms or double-count adjacent years. We interact these event-study indicators, with the variable “$\text{NGO Campaign}_{n,m,y}$,” an indicator variable equal to 1 if firm $n$ experienced an NGO campaign in week $w$ of year $y$, and 0 otherwise. We set the reference month to four months before the AGM, following SEC rules that require the AGM date to be set at least four months in advance.
Figure 2: Firms’ negative media coverage increases if NGOs campaign during AGMs
Note: The figure presents estimates from (1). The dependent variable is an indicator that a firm receives at least one negative E&S news in a week. Standard errors are clustered at the firm level.
@@
The vector “$\mathbf{X}_{n,w,y}$” includes “$\mathbf{1}{\{M_{n,w,y}=\tau\}}$” and “$\text{NGO Campaign}_{n,m,y}$” separately, and controls for (i) the occurrence of negative news and (ii) all E&S news over the previous four weeks to account for a firm’s recent media coverage. [^8] Adding controls for the current count of E&S news to account for NGOs targeting AGMs of high-coverage firms does not qualitatively change the results. Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-week-by-year fixed effects ($\rho_{w,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs). Standard errors are clustered at the firm level.
Figure 2 reports the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients. Negative news follow a bell-shaped pattern around the AGM, peaking significantly above zero in the AGM month, highlighting the benefits of targeting AGMs. [^9] In section 6.2, we show that the spike on AGM date is particularly strong for an NGO’s first few campaigns but it fades away for later ones (Appendix Figure B4). To quantify the magnitude of the AGM peak, an NGO campaign launched in the AGM week doubles the probability of a negative E&S news compared to a campaign initiated four months earlier, on average. After the AGM, coverage returns to baseline levels, indicating that NGOs attract more media attention when campaigning closer to the AGM.
@@
Importantly, we rule out the hypothesis that NGOs select more negative causes when targeting AGMs. Without an adequate ranking of causes by negativity, we estimate (1) with firm $n$’s average E&S news sentiment in week $w$ as the dependent variable. [^10] RavenPack scores the tone of a firm’s news coverage between -1 (adverse) and 1 (favorable). Examples of positive and negative news can be found in Appendix Figure B5. The estimated coefficients show that campaigns launched during AGM months are no different than other campaigns (Appendix Figure B6).
@@ 4.1 Stakeholders’ Reactions
This section explores how campaign timing affects critical stakeholders of a targeted firm, such as shareholders and consumers, while Section 4.2 examines its ramifications for the target’s operations. The relevant exhibits are in the Online Appendix.
@@ 4.1.1 Shareholders
Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022). Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest. We define as “related” a shareholder proposal on the same cause of the campaign in the previous year. Appendix C describes the matching procedure.
@@
Figure B7 shows that only campaigns timed twelve months prior to the next AGM increase the likelihood of a shareholder proposal on the same cause. This substantial increase is as large as 12 percentage points, or 44%, above the average probability of observing a proposal. Thus, NGO campaigns from the previous AGMs likely lead to a shareholder proposal at the current AGM, potentially due to heightened media attention.
@@ 4.1.2 Consumers
Drawing from quarterly data from Compustat, we investigate the responses of consumers to the timing of NGO campaigns by studying the relation between the target’s “Sales to Asset ratio” and the distance between the campaign and the target’s AGM through
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively. The dependent value normalizes $n$’s sales by its asset to account for variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms). Standard errors are clustered at the firm level.
Appendix Figure B8 reports the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients separately for two subsets of the data: for NGO campaigns (i) not planned on an AGM date (blue dots) and (ii) those planned on an AGM date (red triangles). We find that consumers do not respond to NGO campaigns not timed on an AGM day, but they do so to campaigns targeted on the AGM date, implying a loss of sales to total assets of 0.7pp, or about 5%—a magnitude consistent with other papers (Conway and Boxell, 2024; Saint-Jean, 2024).
4.2 Does AGM-Targeting Create Positive Change?
If public shaming were an effective strategy for influencing a firm’s operations, targeting the AGM month should lead to the greatest improvements in a firm’s E&S score. However, Appendix Figure B9 shows that only campaigns launched 3 to 6 months before an AGM result in a subsequent increase in firms’ E&S scores, with a 2 to 3% rise over the following four quarters. In contrast, campaigns launched at other times, including during the AGM month, have an economically and statistically insignificant effect on firms’ scores.
Although ESG scores are an imperfect measure of a firm’s social impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), this finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms. Therefore,
@@ 4.2 Does AGM-Targeting Create Positive Change?
NGO campaigns on an AGM date impact stakeholders but do not lead to improvements in E&S scores. (i) Shareholders are more likely to make related proposals the following year. (ii) The target’s sales drop in the months after a campaign.
5 Timing Campaigns to Influence AGM Votes
Early campaigns could benefit an NGO aiming to influence current AGM outcomes. Since most votes are cast by proxy before the AGM, targeting the AGM date may limit an NGO’s ability to sway a significant share of the votes. For instance, over 85% of vote-eligible shares for Microsoft’s 2023 AGM were cast by proxy before the meeting, indicating that campaigning around AGM dates may have little impact on shareholders’ votes (Microsoft, 2023).
@@ 5 Timing Campaigns to Influence AGM Votes
where “$\text{Success}_{i,n,m,y}$” is an indicator for the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and “$\text{Campaign}_{i,n,m+l,t}$” is an indicator for the occurrence of an NGO campaign on the same cause as proposal $i$, $l$ months before the AGM. The reference period covers the months that are more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$. We account for firm-by-year, and cause-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and causes within firm-year pairs. The standard errors are clustered by firm.
Figure 3: NGOs’ campaigns before the AGM help E&S shareholder proposals succeed
@@
(b) Vote share in favor of proposal
Note: E&S denote environmental and social proposals, respectively. The figure plots the estimated coefficients from (4). The dependent variables are an indicator variable that is one if a shareholder proposal is withdrawn or successful if voted (911 proposals) in Panel (a), and the percentage of votes in favor of a proposal (2,078 proposals) in Panel (b). Standard errors are clustered at the firm level.
Figure 3(a) plots the estimated $\{\hat{\beta}_{l}\}_{l}$. We find that NGO campaigns happening 6 months and 1 month before the AGM positively influence the outcome of related shareholder proposals. They increase, respectively, the probability of success by 2pp (6%) and 3pp (8%) more than a campaign launched more six months earlier.
To understand why campaigns 6 months before the AGM are also somewhat successful, we estimate (4) on the subset of voted proposals, with the percentage of votes in favor as the dependent variable. Note that every proposal to be voted on at the AGM is accompanied by a vote recommendation by the management (“For”, “Against”, or “Abstain”). Institutional investors often follow proxy advisor recommendations, particularly from ISS (Shu, 2024). We include management and ISS recommendations as covariates in (4).
Figure 3(b) shows that only NGO campaigns launched one or three months before the AGM increase support for related shareholder proposals, with an increase of nearly 2 percentage points. Given the generally low baseline support for these proposals (24%), this corresponds to an 8% increase in shareholder support when an NGO campaign addresses the same cause, on average.
In contrast to Figure 3(a), Figure 3(b) detects no significant increase in support for campaigns initiated as early as six (or five) months before the AGM. This suggests that the difference between the two figures arises from shareholder proposals being withdrawn after reaching an agreement with the firm. As a result of these early campaigns, the target firm may take actions to resolve the issue highlighted in both the campaign and the shareholder proposal, leading to an improvement in its E&S score, which supports the finding in Section 4.2 that only early campaigns result in better E&S scores.
Moreover, the fact that campaigns launched one to three months before an AGM help secure more votes aligns with the timing of proxy voting. Shareholders typically receive voting materials within 90 days of the AGM, a period when AGM visibility in the news may increase as well as the salience of the campaign for shareholders (ISS, 2022).
Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015). For this reason, we investigate all E&S proposals in Figure 3(a), which also include withdrawn proposals. Given the low support of proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote. However, those votes can still be helpful in disciplining firms as large shareholder support to a cause may be hard to conceal for managers (Levit and Malenko, 2011).
@@
NGO campaigns are more effective at influencing a firm’s decisions when launched before the AGM.
@@
Sections 4 and 5 outline a tradeoff: late campaigns expose a firm’s wrongdoing to its stakeholders and boost the NGO’s visibility, while early ones are more effective at directly influencing a firm’s decisions. How do NGOs navigate this tradeoff?
@@
6.1 Accumulating Visibility
Targeting AGM-dates could enahnce an NGO’s visibility and its reputation for commitment to a cause. In this section, we explore whether NGOs, after accumulating some visibility, begin launching campaigns earlier—before the target firm’s AGM—to improve their chances of influencing the outcome as suggested in Section 5.
@@ 6.1 Accumulating Visibility
$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}{\{\text{Campaign \#}_{c(\cdot),g,y}=\tau\}}$ (5)
$+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$
which examines “$\text{Distance to AGM}_{c(i,n,g),y}$,” the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using an indicator, “$\text{Campaign \#}_{c,g,y}$,” that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise.
Since our campaign data start in 2010, we do not know how many campaigns an NGO launched before 2010. To control for this truncation as well as time varying confounders such as NGO size, we include NGO-by-year fixed effects. We also include firm-by-year and cause-by-year fixed effects to account for firm and cause-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign. Hence, the identification of $\{\beta_{\tau}\}_{\tau}$ comes from multiple campaigns $c(\cdot)$ by the same NGO $g$ within year $y$ [^11] In Appendix Figure B10, we present similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years. Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample. The standard errors are clustered at the NGO level.
To interpret the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM. Thus, Figure 4 indicates that as the NGO’s experience increases, the new campaigns are launched before the AGM date. [^12] The large negative coefficients estimated for NGOs with many campaigns—e.g., more than 9—are driven by outlier NGOs with several campaigns per year. We postpone the interpretation of this result to Section 7, which sums up and discusses the results.
@@
Alternative mechanism—selection. NGO do not increasingly launch campaigns against firms that may be more easily influenced over time. Appendix Figures B12 and B13 exclude selection on observables. For instance, target sizes do not covary (market capitalization) with NGO experience. In addition, environmental (social) scores do not covary with an NGO’s experience in environmental (social) campaigns.
@@
Figure 4: More experienced NGOs campaign earlier
Note: This figure plots the estimated coefficients from (5). The dependent variable is the months between an NGO campaign and the targeted firm’s AGM date. Only firms with at least two campaigns are considered (1,136 observations). Standard errors are clustered at the NGO level.